According to the May 16,2017 report by Pew Research, 68% of US adults say solar energy is very effective in minimizing air pollution. A sample of 240 U.S adults are surveyed. Let p

ˆ

p^ be the sample proportion of Americans who believe solar energy is very effective in minimizing air pollution. 

  1. What is the population proportion? Answer (decimal form)
  2. What is the sample size? Answer
  3. What is the mean of the sampling proportion? Answer
  4. What is the standard deviation of the sampling proportion? Answer

(Round to 4th decimal place)



1
Expert's answer
2022-05-22T17:38:26-0400

1. p0=0.68p_0=0.68


2. n=240n=240


3. μp^=0.68\mu_{\hat{p}}=0.68


4.

p0(1p0)n=0.68(10.68)2400.0315\sqrt{\dfrac{p_0(1-p_0)}{n}}=\sqrt{\dfrac{0.68(1-0.68)}{240}}\approx0.0315